OpenStudy (anonymous):

expand the brackets 3( 2y + 5) -7(3y-4)

OpenStudy (anonymous):

Distribute 3 and -7 to the respective parenthesis..

OpenStudy (anonymous):

please just give me an answeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeer

OpenStudy (unklerhaukus):

\[3( 2y + 5) -7(3y-4)\] \[(3\times2y + 3\times5) +(-7\times3y+(-7\times-4))\]

OpenStudy (anonymous):

a(b + c) = ab + ac -x(y -z) = -xy + xz

OpenStudy (anonymous):

Don't wanna solve?? You want answer only??